What's the relationship between CO2, bicarbonate (HC03- ), and carbonic acid (H2CO3)? In what form is most CO2 carried in our blood? What is the role of carbonic anhydrase in gas exchange?
Answer #1

The relationship between

Carbon dioxide (CO2) dissolves in with water to form carbonic acid (H2CO3) this reaction is speedup by the enzyme carbonic anhydrase, then carbonic acid dissociates in to bicarbonate (HC03-).

Overall reaction is represented as CO2 + H2O carbonic anhydrase  -à H2CO3 + HCO3- + H+ (this is a reversible reaction and carbonic anhydrase can reverse this reaction – this happens in once RBC reaches lungs)

carbonic acid (H2CO3) and bicarbonate (HC03-) exist in 1:20 in the blood

CO2 / H2CO3 = 1000 / 1

H2CO3 / HCO3- = 1/20

Major form of dissolved CO2 carried in our blood is in the form of bicarbonate (HC03-), 70-75% (rest 25% dissolved in serum or combined with serum protiens and heamoglobin) of the CO2 in our blood is converted in to carbonic acid (H2CO3), which dissociate in to bicarbonate (HC03-) as we know carbonic acid (H2CO3) to bicarbonate (HC03-) ratio is 1:20, that means nearly 71.25% of carbonic acid (H2CO3) gets dissociated in to bicarbonate (HC03-)   

As described above carbonic anhydrase can help H2CO3 + HCO3- formation and reverse reaction of forming CO2 + H2O , In tissues due to high CO2 concentration the reaction is forward (lowers the activation energy of combining C O2 + OH- to give HC03-), whereas in lungs because of low CO2 concentration and high O2 concentration the backward reaction is favored to release CO2 (lowers the activation energy of hydrolyzing HCO3- to CO2 + OH- and release CO2 + H2O)